OpenStudy (anonymous):

use properties of logarithms to find the exact value of the expression, e^log(base(e^2))100 =__________

OpenStudy (mertsj):

\[e ^{\ln e^2}(100)=\]

OpenStudy (mertsj):

Is that it?

OpenStudy (anonymous):

\[e ^{ \left( \log _{e ^{2}}100 \right)}\]

OpenStudy (anonymous):

answer is 10?

OpenStudy (anonymous):

did you just put it in a calc? or how did you get it?

OpenStudy (anonymous):

oke wait please

OpenStudy (anonymous):